Question:
The probabilities that a student will solve Question A and Question B are 0.4 and 0.5 respectively. What is the probability that he solves at least one of the two questions?
एक छात्र द्वारा प्रश्न A और प्रश्न B को हल करने की संभावनाएँ क्रमशः 0.4 और 0.5 हैं। इसकी क्या प्रायिकता है कि वह दोनों में से कम से कम एक प्रश्न हल कर ले?
Options:
- 0.6
- 0.7
- 0.8
- 0.9
- 0.6
- 0.7
- 0.8
- 0.9
Correct Answer: b)
Explanation:
Let events of solving questions A and B be independent.
\(P(A) = 0.4\) and \(P(B) = 0.5\).
The probability of solving at least one question is:
\[P(A \cup B) = P(A) + P(B) – P(A \cap B)\]
Since solving independent questions implies \(P(A \cap B) = P(A) \cdot P(B)\):
\[P(A \cup B) = 0.4 + 0.5 – (0.4 \times 0.5) = 0.9 – 0.20 = 0.70\]
Alternatively: \(1 – P(A’)P(B’) = 1 – (1 – 0.4)(1 – 0.5) = 1 – (0.6)(0.5) = 1 – 0.3 = 0.7\).
Hence, the correct option is (b).
माना प्रश्न A और B हल करने की घटनाएँ स्वतंत्र हैं।
\(P(A) = 0.4\) तथा \(P(B) = 0.5\)।
कम से कम एक प्रश्न हल करने की प्रायिकता:
\[P(A \cup B) = P(A) + P(B) – P(A)P(B)\]
\[P(A \cup B) = 0.4 + 0.5 – (0.4 \times 0.5) = 0.9 – 0.20 = 0.7\]
अतः सही विकल्प (b) है।